OpenStudy (anonymous):

I need an equation for these lines 1) (4,7),(2,3) 2) (9,3),(12,18) 3) (-2,-6),(2,22) 4) (0,0),(5,35) 5) (-3,5),(-9,-1)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

that is a lot of questions !!

OpenStudy (anonymous):

@misty1212 hello

OpenStudy (misty1212):

lets do this one (4,7),(2,3)

OpenStudy (misty1212):

find the slope by \[\frac{y_2-y_2}{x_2-x_1}=\frac{7-3}{4-2}=\frac{4}{2}=2\]

OpenStudy (anonymous):

@misty1212 its a practice worksheet that I have to finish

OpenStudy (misty1212):

then do \[y-y_1=m(x-x_1)=y-3=2(x-2)\]

OpenStudy (misty1212):

solve \[y-3=2(x-2)\]for \(y\) \[y-3=2(x-2)\\ y-3=2x-4\\ y=2x-1\] is your final answer others are similar

OpenStudy (anonymous):

2) y=5x-6 3) y=7x+8